OpenStudy (anonymous):

please help with 4 geometry questions? medal !

OpenStudy (anonymous):

ok ask them...

OpenStudy (anonymous):

@DrPepperx3 ,

OpenStudy (anonymous):

why don't you ask?:)

OpenStudy (anonymous):

sorry my computer keeps freezing

OpenStudy (anonymous):

oh ok...no problem :)

OpenStudy (anonymous):

okay here is question number one(:

OpenStudy (anonymous):

1. What is the length of ? a. 17 b. 15 c. 7 d. 1

OpenStudy (anonymous):

b

OpenStudy (anonymous):

15

OpenStudy (anonymous):

the AD is part of the question and the number line has to do with finding the answer if you could explain how you got the answer it would really help me c:

OpenStudy (anonymous):

count the red lines...

OpenStudy (anonymous):

each of the lines that i draw,mean "+1" for example "-6-7=1" and "-5-6=1" and "1-0=1"

OpenStudy (anonymous):

and the length of the AD is 15x"1" = 15

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

sorry my computer froze again ohh i think i understand now ty(: do you mind helping me with 3 more?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

2. If EF=6 and EG=2, Find the value of FG. The drawing is not to scale a. 17 b. 15 c. 14 d. 6

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

ohhh okay x:

OpenStudy (dumbcow):

how can EG be greater than EF ??

OpenStudy (anonymous):

@dumbcow im not sure thats just what the teacher put on our hw paper im confused :c

OpenStudy (anonymous):

I think the question is quite incorrect :D:)

OpenStudy (anonymous):

@PFEH.1999 i know ! thats what it sounds like lol /:

OpenStudy (dumbcow):

teacher had a typo...it happens teaching is stressful:)

OpenStudy (anonymous):

Do you want to ask other questions?

OpenStudy (anonymous):

@dumbcow yea your right s: well i guess i can just guess lol

OpenStudy (anonymous):

@PFEH.1999 yes(: hold on

OpenStudy (anonymous):

3. What segment is congruent to answe d. none

OpenStudy (anonymous):

AC goes with the question and the number line is to help answer the question s:

OpenStudy (anonymous):

b is correct, BE

OpenStudy (anonymous):

BE = +6 and AC = + 6 too...

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

i think so x: one more question ? and im sorry it takes me awhile to respond my computer is so slow

OpenStudy (anonymous):

no no problem madam :) ask it

OpenStudy (anonymous):

4. If Z is the midpoint of , what are x, RZ, RT? a. x=13 , RZ=38 , and RT=24 b. x=11, RZ=16 , and RT=32 c. x=13 , RZ=24 , and RT=48 d. x=15 , RZ=24 , and RT=48

OpenStudy (anonymous):

where is x?

OpenStudy (anonymous):

@DrPepperx3

OpenStudy (anonymous):

b,c,d are true because 32/16 = 2 , 48/24 = 2 but a is quite incorrect because 24/38 is not 2 but I don't know what is !!!

OpenStudy (anonymous):

i think it wants you to solve for x using RZ and RT

OpenStudy (anonymous):

no in the figure there isn't any x!I think this question is incorrect too :D

OpenStudy (anonymous):

i dont know than /:

OpenStudy (anonymous):

well thank you for all your help(:

OpenStudy (anonymous):

You're welcome , sorry I couldn't help you for all the questions :)